Molecules of a rare gas occur at an average rate of 3 per cubic metre of air and follow a poisson distribution.
i) Find the probability that 3 cubic metre of air contain exactly 4 of the molecules
ii) The probability that at least one molecule is found is 0.99. Find how much air is taken as a sample.
1
Expert's answer
2020-11-05T16:50:29-0500

i. Poisson(3×3)

"P(4)=\frac{9^4e^{-9}}{4!}=0.03374"

ii 1-P(0)=0.99

"1-\frac{(3x)^0e^{-3x}}{0!}=0.99"

"0.01=e^{-3x}"

"\ln0.01=-3x"

"x=\frac{\ln0.01}{-3}=1.5351"

1.5351 cubic metres of air is taken.